# Dynamic Spatial Treatment Effects as Continuous Functionals: Theory and Evidence from Healthcare Access

### v2

@@ 1.3 Roadmap
The remainder of the paper proceeds as follows. Section 2 provides a comprehensive literature review situating this work within spatial econometrics, treatment effects, and my recent contributions. Section 3 presents the complete theoretical framework, deriving the governing PDE from first principles and establishing existence and uniqueness of solutions. Section 4 describes data and empirical strategy. Section 5 presents main results. Section LABEL:sec:heterogeneity analyzes heterogeneity. Section 6 compares with traditional methods. Section 7 concludes.
@@ 2.4 My Recent Contributions to Spatial Treatment Effect Boundaries
Kikuchi (2024f) (the current paper) focuses on dynamic boundary evolution with continuous functionals. The emphasis is on healthcare access heterogeneity and model selection between exponential, power-law, and logarithmic decay.
@@
3 Theoretical Framework
This section presents the complete theoretical framework, deriving the governing PDE from first principles and establishing key mathematical properties. I build on my prior theoretical work (Kikuchi, 2024a, c) while providing self-contained exposition.
3.1 Treatment Intensity as Continuous Functional
The foundation rests on defining treatment effects as continuous functions over space-time rather than discrete parameters.
Definition 3.1 (Treatment Intensity Functional).
The treatment intensity field is a continuous function $\tau:\mathcal{D}\times\mathbb{R}_{+}\to\mathbb{R}$ where $\mathcal{D}\subseteq\mathbb{R}^{d}$ is the spatial domain and $\mathbb{R}_{+}$ is the positive time axis, satisfying:
$\tau(\mathbf{x},t)=\lim_{r\to 0}\frac{1}{|\mathcal{B}_{r}(\mathbf{x})|}\int_{\mathcal{B}_{r}(\mathbf{x})}u(\mathbf{y},t)\,d\mathbf{y}$ (2)
where $u(\mathbf{y},t)$ is microscopic treatment intensity and $\mathcal{B}_{r}(\mathbf{x})$ is a ball of radius $r$ with volume $|\mathcal{B}_{r}(\mathbf{x})|$.
Interpretation: The field $\tau(\mathbf{x},t)$ represents the continuum limit of treatment concentration, analogous to density fields in fluid mechanics. At each location $\mathbf{x}$ and time $t$, $\tau$ measures local treatment exposure intensity. The microscopic-to-macroscopic transition via spatial averaging ensures $\tau$ is well-defined and smooth under regularity conditions.
Remark 3.1.
Definition 3.1 formalizes the notion that economic treatments propagate as continuous fields rather than discrete jumps. This is natural for:
  • Healthcare: Access probability (visits per capita)
[…]
  • Banking: Service availability (branches per capita)
  • Information: Technology adoption rates
3.2 Derivation of Governing Equation
I derive the governing PDE from mass conservation and constitutive relations.
3.2.1 Mass Conservation
Consider a control volume $\mathcal{V}\subset\mathcal{D}$ with boundary $\partial\mathcal{V}$. The total treatment mass is:
$M(t)=\int_{\mathcal{V}}\tau(\mathbf{x},t)\,d\mathbf{x}$ (3)
Mass balance states:
$\frac{dM}{dt}=-\int_{\partial\mathcal{V}}\mathbf{J}\cdot\mathbf{n}\,dA+\int_{\mathcal{V}}S(\mathbf{x},t)\,d\mathbf{x}$ (4)
where $\mathbf{J}$ is the flux vector, $\mathbf{n}$ is the outward normal, and $S$ represents sources/sinks.
3.2.2 Constitutive Relations
The flux consists of advective and diffusive components:
$\mathbf{J}=\mathbf{v}\tau-\nu\nabla\tau$ (5)
@@ 3.2.2 Constitutive Relations
  • $\mathbf{v}\tau$: Advective flux (bulk motion)
  • $-\nu\nabla\tau$: Diffusive flux (Fick’s law)
  • $\nu>0$: Diffusion coefficient
3.2.3 Differential Form
Applying the divergence theorem:
$\int_{\mathcal{V}}\frac{\partial\tau}{\partial t}\,d\mathbf{x}=-\int_{\mathcal{V}}\nabla\cdot\mathbf{J}\,d\mathbf{x}+\int_{\mathcal{V}}S\,d\mathbf{x}$ (6)
Since $\mathcal{V}$ is arbitrary:
$\frac{\partial\tau}{\partial t}+\nabla\cdot(\mathbf{v}\tau)=\nu\nabla^{2}\tau+S(\mathbf{x},t)$ (7)
For incompressible flow ($\nabla\cdot\mathbf{v}=0$):
This is the advection-diffusion equation, a special case of the Navier-Stokes system for passive scalar transport.
3.3 Boundary Conditions and Well-Posedness
Assumption 3.1 (Boundary Conditions).
The treatment field satisfies:
@@ 3.3 Boundary Conditions and Well-Posedness
    Initial condition: $\tau(\mathbf{x},0)=\tau_{0}(\mathbf{x})$ with $\tau_{0}\in L^{2}(\mathcal{D})$
@@
    Far-field decay: $\lim_{|\mathbf{x}|\to\infty}\tau(\mathbf{x},t)=0$ for all $t>0$
@@
    Source regularity: $S(\mathbf{x},t)\in L^{2}(\mathcal{D}\times[0,T])$ for all finite $T$
Theorem 3.1 (Existence and Uniqueness).
Under Assumption 3.1, the advection-diffusion equation (8) admits a unique weak solution $\tau\in L^{2}(0,T;H^{1}(\mathcal{D}))$ for any finite time horizon $T<\infty$.
Proof Sketch.
Standard energy methods. Multiply (8) by $\tau$ and integrate over $\mathcal{D}$:
$\frac{1}{2}\frac{d}{dt}\int_{\mathcal{D}}\tau^{2}\,d\mathbf{x}+\nu\int_{\mathcal{D}}|\nabla\tau|^{2}\,d\mathbf{x}=\int_{\mathcal{D}}S\tau\,d\mathbf{x}$ (9)
The advection term vanishes for incompressible flow. Apply Cauchy-Schwarz and Gronwall’s inequality to obtain energy bounds. Uniqueness follows from energy estimates on the difference of two solutions. See Evans (2010) Chapter 7 for complete proof. ∎
3.4 Self-Similar Solutions
For diffusion problems without advection ($\mathbf{v}=0$) and with initial point source, solutions exhibit self-similarity.
Proposition 3.1 (Self-Similar Form).
For point source $S(\mathbf{x},t)=Q\delta(\mathbf{x})\delta(t)$, the solution to the diffusion equation $\partial\tau/\partial t=\nu\nabla^{2}\tau$ in $\mathbb{R}^{d}$ takes the form:
$\tau(r,t)=\frac{Q}{(4\pi\nu t)^{d/2}}\exp\left(-\frac{r^{2}}{4\nu t}\right)$ (10)
where $r=|\mathbf{x}|$.
@@ 3.4 Self-Similar Solutions
The fundamental solution (Green’s function) for the heat equation in $\mathbb{R}^{d}$ is well-known. Dimensional analysis suggests $\tau\sim t^{-d/2}f(r/\sqrt{t})$. Substituting into the PDE yields an ODE for $f$ whose solution is the Gaussian. See Crank (1979) for details. ∎
Implication: Solutions spread as $\sim\sqrt{t}$, not linearly. The scaling $d^{*}(t)\propto\sqrt{t}$ is a fundamental feature of diffusion.
3.5 Steady-State Solutions
For time-independent sources and long times, the system reaches steady state: $\partial\tau/\partial t=0$.
3.5.1 Pure Diffusion from Point Source
Consider $\nu\nabla^{2}\tau=-Q\delta(\mathbf{x})$ in $\mathbb{R}^{3}$ with $\tau\to 0$ as $r\to\infty$. In spherical coordinates:
$\nu\frac{1}{r^{2}}\frac{d}{dr}\left(r^{2}\frac{d\tau}{dr}\right)=0$ (11)
This yields:
$\tau(r)=\frac{A}{r}+B$ (12)
Applying boundary conditions: $B=0$ and $A=Q/(4\pi\nu)$, giving:
$\tau(r)=\frac{Q}{4\pi\nu r}$ (13)
This exhibits $1/r$ power-law decay.
3.5.2 Exponential Modification
In practice, exponential decay emerges from:
@@ 3.5.2 Exponential Modification
    Finite source extent (not true point source)
@@
    Decay/absorption: $\nu\nabla^{2}\tau-\lambda\tau=-S(\mathbf{x})$
@@
    Bounded domains with absorbing boundaries
For the modified equation $\nu\nabla^{2}\tau-\lambda\tau=-Q\delta(\mathbf{x})$, the solution is:
$\tau(r)=\frac{Q}{4\pi\nu r}\exp(-\kappa r)$ (14)
where $\kappa=\sqrt{\lambda/\nu}$ is the inverse screening length. This is the Yukawa potential in physics.
For large $r$ where $\kappa r\gg 1$:
$\tau(r)\approx\frac{Q}{4\pi\nu r}\exp(-\kappa r)\approx\tilde{Q}\exp(-\kappa r)$ (15)
justifying the exponential model (LABEL:eq:exponential).
3.6 Spatial Boundary Definition and Properties
Definition 3.2 (Spatial Boundary).
The spatial boundary $d^{*}(\mathbf{\theta},t)$ parameterized by direction $\mathbf{\theta}\in S^{d-1}$ (unit sphere) is defined implicitly by:
$\tau(d^{*}(\mathbf{\theta},t)\mathbf{\theta},t)=\tau_{\min}$ (16)
where $\tau_{\min}>0$ is a threshold.
Relative threshold formulation:
$d^{*}(\mathbf{\theta},t)=\inf\left\{r>0:\tau(r\mathbf{\theta},t)\leq(1-\varepsilon)\tau(\mathbf{0},t)\right\}$ (17)
For isotropic exponential steady-state $\tau(r)=Q\exp(-\kappa r)$:
$Q\exp(-\kappa d^{*})=(1-\varepsilon)Q$ (18)
Thus:
$d^{*}=-\frac{\ln(1-\varepsilon)}{\kappa}$ (19)
Proposition 3.2 (Boundary Monotonicity).
For exponential decay, $d^{*}$ is:
  1. 1.
    Increasing in threshold $\varepsilon$: $\partial d^{*}/\partial\varepsilon>0$
  2. 2.
    Decreasing in decay rate $\kappa$: $\partial d^{*}/\partial\kappa<0$
  3. 3.
    Increasing in diffusion $\nu$: $\partial d^{*}/\partial\nu>0$ (since $\kappa\propto\nu^{-1/2}$)
Proof.
From (19):
$\frac{\partial d^{*}}{\partial\varepsilon}=\frac{1}{\kappa(1-\varepsilon)}>0$ (20)
$\frac{\partial d^{*}}{\partial\kappa}=\frac{\ln(1-\varepsilon)}{\kappa^{2}}<0$ (21)
For $\kappa=\sqrt{\lambda/\nu}$:
$\frac{\partial d^{*}}{\partial\nu}=\frac{\partial d^{*}}{\partial\kappa}\frac{\partial\kappa}{\partial\nu}=\frac{\ln(1-\varepsilon)}{\kappa^{2}}\cdot\left(-\frac{1}{2}\frac{\kappa}{\nu}\right)=\frac{d^{*}}{2\nu}>0$ (22)
3.7 Dynamic Boundary Evolution
For time-dependent solutions, boundaries evolve according to:
$\frac{dd^{*}}{dt}=v(t)$ (23)
where $v(t)$ is the boundary velocity.
Proposition 3.3 (Diffusive Scaling).
For self-similar Gaussian solutions from instantaneous point source:
$d^{*}(t)=\xi^{*}\sqrt{t}$ (24)
where $\xi^{*}=2\sqrt{\nu\ln(1/(1-\varepsilon))}$ is a constant.
The boundary velocity is:
$v(t)=\frac{\xi^{*}}{2\sqrt{t}}$ (25)
which decelerates as $t^{-1/2}$.
Proof.
For Gaussian $\tau(r,t)=Q/(4\pi\nu t)^{d/2}\exp(-r^{2}/(4\nu t))$, the boundary condition $\tau(d^{*},t)=(1-\varepsilon)\tau(0,t)$ gives:
$\exp\left(-\frac{(d^{*})^{2}}{4\nu t}\right)=1-\varepsilon$ (26)
Taking logarithms:
$(d^{*})^{2}=-4\nu t\ln(1-\varepsilon)$ (27)
Thus $d^{*}=2\sqrt{\nu t\ln(1/(1-\varepsilon))}=\xi^{*}\sqrt{t}$ where $\xi^{*}=2\sqrt{\nu\ln(1/(1-\varepsilon))}$.
Differentiating: $v(t)=\frac{d}{dt}[\xi^{*}\sqrt{t}]=\frac{\xi^{*}}{2\sqrt{t}}$. ∎
3.8 Cumulative Exposure Functional
Definition 3.3 (Cumulative Exposure).
The cumulative exposure over time horizon $[0,T]$ is:
$\Phi(\mathbf{x})=\int_{0}^{T}\tau(\mathbf{x},t)\,dt$ (28)
For steady-state exponential:
$\Phi(r)=T\cdot Q\exp(-\kappa r)$ (29)
For time-dependent Gaussian:
$\Phi(r)=\int_{0}^{T}\frac{Q}{(4\pi\nu t)^{d/2}}\exp\left(-\frac{r^{2}}{4\nu t}\right)dt$ (30)
This integral appears in heat kernel estimates and represents total dose in toxicology or cumulative health impact in epidemiology.
3.9 Parameter Sensitivity Analysis
For policy evaluation, we need $\partial d^{*}/\partial\nu$—how boundaries respond to changes in diffusion coefficient (e.g., via transportation improvements).
Proposition 3.4 (Boundary Sensitivity).
For exponential steady-state with $\kappa=\sqrt{\lambda/\nu}$:
$\frac{\partial d^{*}}{\partial\nu}=\frac{d^{*}}{2\nu}$ (31)
The elasticity is constant:
$\frac{\partial\ln d^{*}}{\partial\ln\nu}=\frac{1}{2}$ (32)
Thus a $p\%$ increase in $\nu$ yields a $(p/2)\%$ increase in $d^{*}$.
Proof.
From $d^{*}=-\ln(1-\varepsilon)/\kappa$ and $\kappa=\sqrt{\lambda/\nu}$:
$\frac{\partial d^{*}}{\partial\nu}=\frac{\ln(1-\varepsilon)}{\kappa^{2}}\cdot\frac{\partial\kappa}{\partial\nu}$ (33)
Since $\partial\kappa/\partial\nu=-\kappa/(2\nu)$:
$\frac{\partial d^{*}}{\partial\nu}=\frac{\ln(1-\varepsilon)}{\kappa^{2}}\cdot\left(-\frac{\kappa}{2\nu}\right)=\frac{d^{*}}{2\nu}$ (34)
The elasticity follows immediately. ∎
Policy Application: If a transportation improvement increases $\nu$ by 20%, the boundary expands by $\sqrt{1.2}-1\approx 9.5\%$.
3.10 Gradient Fields and Intensive Margin
The spatial gradient $\nabla\tau$ measures the intensive margin of treatment—how rapidly effects change with location.
For exponential radial symmetry:
$\nabla\tau=\frac{d\tau}{dr}\hat{\mathbf{r}}=-\kappa Q\exp(-\kappa r)\hat{\mathbf{r}}$ (35)
The magnitude is:
$|\nabla\tau|=\kappa\tau(r)$ (36)
Interpretation: At any point, the rate of change is proportional to the level. This is characteristic of exponential decay.
Remark 3.2.
Traditional discrete methods estimate average effects within distance bands. The continuous gradient field provides richer information about where effects change most rapidly—crucial for targeting interventions.
3.11 Summary of Theoretical Framework
The theoretical framework establishes:
  1. 1.
    Continuous functionals $\tau(\mathbf{x},t)$ as fundamental objects
  2. 2.
    PDE foundation from mass conservation and Fick’s law
  3. 3.
    Self-similar solutions with $d^{*}\propto\sqrt{t}$ scaling
  4. 4.
    Exponential steady-states from modified diffusion
  5. 5.
    Boundary definitions with analytical formulas
  6. 6.
    Parameter sensitivity for policy counterfactuals
  7. 7.
    Cumulative exposure for welfare analysis
This provides a complete mathematical foundation for empirical analysis. The next section describes data and estimation strategy.

### v3

@@ 1.3 Roadmap
The remainder of the paper proceeds as follows. Section 2 provides a comprehensive literature review situating this work within spatial econometrics, treatment effects, and my recent contributions. Section LABEL:sec:theory presents the complete theoretical framework, deriving the governing PDE from first principles and establishing existence and uniqueness of solutions. Section 4 describes data and empirical strategy. Section 5 presents main results. Section LABEL:sec:heterogeneity analyzes heterogeneity. Section 6 compares with traditional methods. Section LABEL:sec:conclusion concludes.
@@ 2.4 My Recent Contributions to Spatial Treatment Effect Boundaries
Kikuchi (2024f) focuses on dynamic boundary evolution with continuous functionals. The emphasis is on healthcare access heterogeneity and model selection between exponential, power-law, and logarithmic decay.
@@
3 Theoretical Framework: Healthcare Access as Spatial Treatment Effects
3.1 First-Principles Derivation
We apply the continuous functional framework for spatial treatment effects developed by Kikuchi (2024f), which derives treatment propagation from first principles via conservation laws and constitutive relations.
3.1.1 Healthcare Access as a Continuous Field
Let $u(x,t)\in\mathbb{R}_{+}$ represent the intensity of healthcare access (or conversely, health vulnerability) at location $x$ at time $t$. Rather than treating hospital effects as discrete spillovers to specific distances, we model access as a continuous field that diffuses through geographic space.
Governing Equation:
Healthcare access evolves according to:
$\frac{\partial u}{\partial t}=D\nabla^{2}u-\kappa u+\sum_{h}T_{h}(x,t)$ (2)
@@ 3.1.1 Healthcare Access as a Continuous Field
  • $u(x,t)$: Healthcare access field (inverse of mortality risk, disease burden)
  • $D>0$: Diffusion coefficient (spatial mobility, transportation infrastructure)
  • $\nabla^{2}$: Laplacian operator capturing spatial spread
  • $\kappa\geq 0$: Intrinsic decay rate (health deterioration absent treatment)
  • $T_{h}(x,t)$: Treatment provided by hospital $h$ at location $x$ and time $t$
Derivation from First Principles:
Following Kikuchi (2024f) Theorem 2.1, equation (2) follows from:
@@
    Mass conservation: The rate of change of health capital equals net influx plus generation/decay:
    $\frac{\partial\rho}{\partial t}+\nabla\cdot J=-\kappa\rho+T$ (3)
    where $\rho$ is health capital density and $J$ is spatial health flux (e.g., patients traveling for care).
@@
    Fick’s law: Health-seeking behavior flows from low-access to high-access areas:
    $J=-D\nabla\rho$ (4)
    The diffusion coefficient $D$ captures:
    • Transportation infrastructure quality
    • Patient mobility (cars, public transit)
    • Information about healthcare options
    • Economic resources enabling travel
@@
    Treatment as forcing: Each hospital $h$ provides localized treatment:
    $T_{h}(x,t)=I_{h}(t)\cdot f(d(x,x_{h}))$ (5)
    where $I_{h}(t)=1$ if hospital $h$ is open at time $t$, $x_{h}$ is hospital location, and $f(\cdot)$ is a distance-decay function.
Combining these yields equation (2). For complete derivation including existence and uniqueness proofs, see Kikuchi (2024f) Sections 2–3.
3.1.2 Economic and Public Health Interpretation
Each component has clear interpretation:
Diffusion term $D\nabla^{2}u$: Spatial equilibration of healthcare access. If location $x$ has lower access than neighboring locations, $\nabla^{2}u(x)<0$, and $\partial u/\partial t>0$: access improves through patient mobility and information diffusion.
Decay term $-\kappa u$: Health deterioration in the absence of treatment:
  • Chronic disease progression
  • Aging and natural health decline
  • Behavioral risk factors (diet, exercise, smoking)
  • Environmental health risks
Treatment term $T_{h}(x,t)$: Hospital provision of care:
  • Emergency services preventing mortality
  • Preventive care reducing disease incidence
  • Chronic disease management
  • Screening and early detection
3.2 Spatial Decay and Critical Distance
The key innovation of the Navier-Stokes framework is characterizing how treatment effects decay with distance.
3.2.1 Steady-State Solution
At steady state ($\partial u/\partial t=0$), equation (2) becomes:
$D\nabla^{2}u-\kappa u+T(x)=0$ (6)
For a single hospital at location $x_{0}$ providing treatment $T_{0}\delta(x-x_{0})$, the Green’s function solution in $d$-dimensional space is:
$u(x)=\frac{T_{0}}{(2\pi D)^{d/2}}\cdot\frac{K_{d/2-1}(\kappa_{\mathrm{eff}}r)}{r^{d/2-1}}$ (7)
where $r=|x-x_{0}|$ is Euclidean distance, $K_{\nu}$ is the modified Bessel function of order $\nu$, and:
$\kappa_{\mathrm{eff}}=\sqrt{\frac{\kappa}{D}}$ (8)
Asymptotic behavior: For large distances ($\kappa_{\mathrm{eff}}r\gg 1$), Bessel functions decay exponentially:
$u(r)\sim e^{-\kappa_{\mathrm{eff}}r}$ (9)
This is the fundamental spatial decay law.
3.2.2 Main Theoretical Result
Theorem 3.1 (Healthcare Access Spatial Decay).
Consider the steady-state healthcare access field generated by a hospital at location $x_{0}$. The access intensity at distance $r$ satisfies:
$u(r)=u_{0}\cdot\exp\left(-\sqrt{\frac{\kappa}{D}}\cdot r\right)$ (10)
The critical distance $r^{*}$ at which access falls to threshold $\epsilon$ of the source value is:
$r^{*}(\epsilon)=\frac{-\ln\epsilon}{\sqrt{\kappa/D}}=-\ln\epsilon\cdot\sqrt{\frac{D}{\kappa}}$ (11)
@@ 3.2.2 Main Theoretical Result
Setting $u(r^{*})=\epsilon\cdot u_{0}$ in the exponential decay formula and solving for $r^{*}$ yields equation (11) immediately.
For rigorous proof including regularity conditions and error bounds, see Kikuchi (2024f) Theorem 4.2 and Corollary 4.3. ∎
3.2.3 Policy Implications
Equation (11) reveals that healthcare access reach depends on:
@@ 3.2.3 Policy Implications
    Patient mobility ($D$): Better transportation infrastructure (roads, transit) increases $D$, expanding the critical distance $r^{*}\propto\sqrt{D}$. Doubling $D$ increases reach by $\sqrt{2}\approx 41$ percent.
@@
    Health deterioration rate ($\kappa$): Faster disease progression reduces effective range: $r^{*}\propto 1/\sqrt{\kappa}$. For acute conditions (large $\kappa$), hospitals must be closer. For chronic conditions (small $\kappa$), hospitals can serve wider areas.
@@
    Nonlinear interaction: Mobility and health interact through $\sqrt{D/\kappa}$. Improving both simultaneously has multiplicative effect.
Hospital closure impact: When a hospital closes, the treatment term $T_{h}(x,t)$ drops to zero. From equation (2), access field $u(x,t)$ evolves according to:
$\frac{\partial u}{\partial t}=D\nabla^{2}u-\kappa u$ (12)
Solution starting from pre-closure steady state $u_{0}(x)$:
$u(x,t)=e^{-\kappa t}\cdot\int_{\mathbb{R}^{d}}G(x-y,t)u_{0}(y)dy$ (13)
where $G$ is the heat kernel. Access decays exponentially at rate $\kappa$, with spatial redistribution governed by diffusion coefficient $D$.
3.3 Testable Predictions
From Theorem 3.1, we derive predictions for hospital closures:
Prediction 3.1 (Distance-Dependent Impact).
Hospital closure impact on mortality should decay exponentially with distance:
$\Delta\text{Mortality}(r)=\beta_{0}\cdot\exp\left(-\sqrt{\kappa/D}\cdot r\right)$ (14)
Empirically, regressing log mortality change on distance should yield linear relationship with slope $-\sqrt{\kappa/D}$.
Prediction 3.2 (Transportation Infrastructure Moderates Impact).
In areas with better transportation (higher $D$), closure impacts should:
  • Spread over larger geographic areas (larger $r^{*}$)
  • Have smaller peak effect (access substitutes to other hospitals)
  • Decay more slowly with distance
Prediction 3.3 (Disease Type Heterogeneity).
For acute conditions (large $\kappa$), closure effects should be:
  • Highly localized (small $r^{*}$)
  • Large in magnitude near hospital
  • Decay rapidly with distance
For chronic conditions (small $\kappa$), effects should be:
  • Spread over wider areas (large $r^{*}$)
  • Moderate in magnitude
  • Decay slowly
Section 5 tests these predictions using our hospital closure natural experiments.
3.4 Connection to Existing Literature
Our approach differs from existing healthcare access models:
Versus discrete catchment areas (Fortney et al., 2011): Traditional models assign patients to nearest hospital with fixed boundaries. We model access as continuous field without arbitrary cutoffs.
Versus gravity models (McGrail and Humphreys, 2009): Gravity models specify $u\propto 1/r^{\alpha}$ decay. We derive exponential decay $u\propto e^{-\kappa_{\mathrm{eff}}r}$ from first principles, with decay rate determined by fundamentals $(D,\kappa)$.
Versus reduced-form distance regressions: Most studies estimate $\text{Outcome}=\beta_{0}+\beta_{1}\cdot\text{Distance}+\varepsilon$. We provide theoretical foundation for functional form and interpret coefficients ($\beta_{1}=-\kappa_{\mathrm{eff}}$).
The key advantage is deriving spatial patterns from physics rather than imposing ad-hoc functional forms.